Question 1 0/1 View Policies Show Attempt History Current Attempt in Progress . Your answer is incorrect. Thomas Taylor justShow work please.
Answer #1

Information provided:

Par value= future value= $1,000

Time= 5 years

Coupon rate= 5.5%

Coupon payment= 0.055*1,000= $55

Market interest rate= 10%

The price of the bond is calculated by computing the present value.

Enter the below in a financial calculator to compute the present value:

FV= 1,000

N= 5

PMT= 55

I/Y= 10

Press the CPT key and PV to compute the present value.

The value obtained is 829.41.

Therefore, the maximum amount that Thomas should be willing to pay for the bond is $829.41.

In case of any query, kindly comment on the solution.